Locate the absolute maximum and minimum for each of the following functions and justify your responses. (Show all work)


1) 𝑓(π‘₯) = square root of x^2 βˆ’ π‘₯ on the interval [βˆ’1,1]

2) 𝑔(π‘₯) = π‘₯𝑒^2π‘₯ on the interval [βˆ’2,0]


1
Expert's answer
2021-02-09T03:16:49-0500

"Solution: (i)~Given ~ that ~: f(x)=\sqrt {x^2-x } ~on ~[-1,1]
\\We~find ~ f'(x)=(2x-1) \frac{1}{2(x^2 -x )^{\frac{1}{2}}}
\\Now ~solving ~ for ~ (2x-1) \frac{1}{2(x^2 -x )^{\frac{1}{2}}}=0
\\we ~get~ x=\frac{1}{2}
\\ \therefore ~we~ find~ f(x)~ on~~ the ~ points x=-1,\frac{1}{2},1
\\ \therefore we~ get,f(-1)= \sqrt{2}=1.41421
\\~~~~~~~~~~~~~~~~~~~~f(\frac{1}{2})=0.8660~and
\\~~~~~~~~~~~~~~~~~~~~~f(0)=0
\\ \therefore (2x-1) \frac{1}{2(x^2 -x )^{\frac{1}{2}}}
\\\therefore Absolute~ Maximum~=1.41421~ and ~Absolute ~minimum~=~0
\\(ii)~g(x)=xe^{2x}~on~[-2,0]
\\ \therefore g'(x)=2xe^{2x}+e^{2x}=e^{2x}(2x+1)
\\ Solving, e^{2x}(2x+1)=0, We~ get,
\\x=-\frac{1}{2}
\\ \therefore we~find ~ f(x)~at ~points~x=-2,-\frac{1}{2},0
\\ \therefore f(-2)=-0.036631
\\f(-\frac{1}{2})=-\frac{1}{2e}=-0.183939720~and ~
\\f(0)=0
\\\therefore Absolute~ Maximum~=0~ and ~Absolute ~minimum~=~-0.183939720"